Suppose you have a 12.0V motorcycle battery that can move 5000C of charge and a 12.0V car battery that can move 60, 000C of charge. how much energy does each deliver? (Assume that the numerical value of each charge is accurate to three significant figures)

1
Expert's answer
2022-03-03T12:04:55-0500
"W_1=\frac{q_1V}{2}=\frac{5000*12.0}{2}=30000\:\rm J"

"W_2=\frac{q_2V}{2}=\frac{60000*12.0}{2}=360000\:\rm J"